OpenStudy (anonymous):

What is the exact value for the expression the square root of 52. − the square root of 13. + the square root of 117.? Simplify if possible. A:4 the square root of 13. B:the square root of 13. C:2 the square root of 39. D:8 the square root of 39.

OpenStudy (anonymous):

\[\Large \sqrt{52}-\sqrt{13}+\sqrt{117}\] What are the factors of each number? Hint: 13 is a factor of each one.

OpenStudy (anonymous):